hi I had this question on a test and i just couldn't figure out the answer. the test was on the elimination method.
heres the problem: 5x+y=1, 10x-2y=7
please help, thank you!

Answer by checkley77(12569)

5x+y=1 multiply by 2 & add.
10x-2y=7
10x+2y=2
--------------------
20x=9
x=9/20 ans.
5*9/20+y=1
45/20+y=1
y=1-45/20
y=20/20-45/20
y=-25/20
y=-5/4 ans.
Proof:
10*9/20-2*-5/4=7
90/20+10/4=7
4.5+2.5=7
7=7